ASHANTI AND MARIA WENT TO THE STORE TO BUY SNACKS FOR THEIR BACK TO SCHOOL PARTY. THEY BOUGHT BAGS OF CHIPS,PRETZELS,NACHOS. THEY BOUGHT FOUR TIMES AS MANY BAGS OF PRETZELS AS BAGS OF CHIPS, AND THREE FEWER BAGS OF NACHOS THAN BAGS OF PRETZELS. IF x REPRESENTS THE NUMBER OF BAGS OF CHIPS THEY BROUGHT, EXPRESS, IN TERMS OF x, HOW MANY BAGS OF SNACKS THEY BROUGHT IN ALL
Answer by ankor@dixie-net.com(22740)


:
Three types of junk food: x(chips, p(pretzels), n(nachos)
:
Four times as many bags of p as x:
p = 4x
:
Three fewer bags of n as p
n = p - 3
:
Substitute 4x for p
n = 4x - 3
:
Total bags = x + p + n
:
f(x) = x + 4x + (4x - 3)
f(x) = x + 4x + 4x - 3
f(x) = 9x - 3